lifewmylilweeb:

What is the solution set of y = x2 + 2x + 7 and y = x + 7? A. {(0, 7), (-1, 6)} B. {(0, 7), (-7, 0)} C. {(0, 7), (1, 8)} D. {(-2, 0), (4, 0)}

surjithayer:

any idea?

lifewmylilweeb:

i think its A

GeekGold:

Is that x^2?

lifewmylilweeb:

yes

surjithayer:

soltion is the point where two equations have a common point. or both equations are equal. solving \[x^2+2x+7=x+7\] or \[x^2+2x+7-x-7=0\] \[x^2+x=0\] \[x(x+1)=0\] x=?

GeekGold:

Just put both equations into a graphing calculator and find where the intersect, that's the answers